0-100-s hop de 1 There is The device
show formulas using excel
cell refrences
Answer #1

Using the data, the demand curve can be drawn as follows.

Price 20 40 60 80 100 Quantity

Point elasticity of demand can be calculated as e = bad Now given a demand function Q = 100 – 5P, E= (-5). Using this formula

The first table shows the calculated values of the elasticity at each of the price-quantity combinations and the second table shows the formula view of the calculations. Note that the elasticity numbers are rounded up to 2 decimal places.

B 0 C Elasticity -0.05 -0.11 -0.18 -0.25 А 1P 21 3 2 4 3 5 4 65 00 10 9 11 10 12 11 13 12 =100-5 A2 -100-5 A3 =100-5 A4 |=100

Now using the calculated elasticity, we can answer the remaining question. We know that when absolute value of elasticity is greater than 1, the demand is elastic, when it is less than 1, demand is inelastic and when it is equal to 1, the demand has unitary elasticity.

The demand curve is inelastic at price from 1 to 9.

The demand curve has unitary elasticity at price 10

The demand curve is elastic at price from 11 to 19